A traveler left a town A at 5:00 am, and traveled toward a town B. At 7:00 o’clock another traveler left town B and traveled toward A. They met at 9:05 am. The first traveler arrived at B and the second at A at the same time. How long did each make the journey?
Answer by Edwin McCravy(20054)

A traveler left a town A at 5:00 am, and traveled toward a town B. At 7:00
o’clock another traveler left town B and traveled toward A. They met at 9:05 am.
The first traveler arrived at B and the second at A at the same time. How long
did each make the journey?

Let traveler 1's rate be r
Let traveler 2's rate be s
Suppose they each arrive at their destinations at t hours after
meeting at M

From 5:00AM to 7:00AM is 2 hours
From 7:00AM to 9:05AM is 2 hours 5 minute = 1 1/12 hours = 25/12 hours

A------P---------M-------B

Suppose traveler 1 is at P at 7:00 when traveler 2 starts.

Suppose they meet at M.

Using distance = rate*time for traveler 1:

Distance AP = 2r
Distance PM = (25/12)r
Adding, distance AM = AP+PM = 2r + (25/12)r = (49/12)r
Distance MB = tr

For traveler 2

Distance BM = (25/12)s
Distance MA = ts

--------------------------------

BM = MB, so (25/12)s = tr

or

(1)   25s = 12tr

MA = AM, so ts = (49/12)r

or

(2)  12ts = 49r

Dividing equals by equals, from (1) and (2) we get

25/(12t) = (12t)/49

144t^2 = 1225

   t^2 = 1225/144

     t = 35/12 = 2 11/12 hours = 2 hours and 55 minutes.

After meeting at 9:05AM they arrived at their respective towns at 12 noon.

So traveler 1's trip took 7 hours and traveler 2's trip took 5 hours.

Edwin